stuelenT pexomel tha eachons that we perfomed in-the termodynnvnics lab, And obtained the dA#4 shown n the SbLa
ka renchms woRe all nn in 100.o,n1イHoo, starting at a Mgo ka wafe-be rf ,yo^ used 0.015 2 moles oumase the readinos weee mstud nn in loo.0mlbf ethanol
stuelenT pexomel tha eachons that we perfomed in-the termodynnvnics lab, And obtained the dA#4 shown n the SbLa
Answer #1

#1: Moles of MgO taken = 0.0152 mol

Heat released by 0.0152 mol MgO = 128 kJ/mol * 0.0152 mol = 1.9456 kJ *(1000 J/1kJ) = 1945.6 J

Since the reaction is exothermic, 1945.6 J heat will be used in increasing the temperature of water.

Given mass of water, m = 100.0 mL * (1g/1mL) = 100.0 g

Initial temperature, Ti = 22.4 oC.

Let the final temeprature be T

=> Q = 1945.6 J = m*s*\DeltaT

=> 1945.6 J = 100.0 * 4.184 J/g,oC * (T - 22.4)

=> T - 22.4 = 4.65

=> T = 4.65 + 22.4 = 27.1 oC (Answer)

#2: The specific heat capacity of ethanol is 2.57 J/g.oC which is less than that of water. Due to lesser specific heat capacity of ethanol, heat absorbed by ethanol is lesser. Hence the final temperature will be lesser than that of water.